### v1

@@ V.2 Emulation in ($\mathrm{Re}\,E$, $\mathrm{Im}\,E$) with fixed $H$ and sources
To check this explanation, we apply a (near)-optimal rational approximant to analytically continue the same training data. The approximant takes the form of Eq. (127). The BCT poles are exponentially distributed according to Eq. (128), but we explore different $\theta_{\text{BCT}}$ angles. The branch point location, as needed for placing the BCTs, is taken to be $E=0$. As for the bound state pole locations needed in Eq. (127), we use the $E_{B}$ result from a given emulator and plug $\mathrm{Re}\,E_{B}$ as the bound state location in the approximant—$\mathrm{Im}\,E_{B}$ is tiny. Eventually, the free parameters, including the pole residues and the location of the resonance $E_{R}$, are fitted against the training data $\mathcal{A}(E^{tr}_{\alpha})$.
Fig. 19 shows a set of results with $\theta_{\text{BCT}}=0.75\pi$. In the first two panels, due to the different horizontal scales on the left and right sides of the plots, the discretized BCTs seem bent, but they are on straight lines. With this extra step and a large $\theta_{\text{BCT}}$, the resonances, extracted based on the training data and partial information from the previous emulators, now agree with the exact result. This also shows that the resonance can be identified as an isolated pole and, thus, part of the eigenstates. The 3rd and the 4th panels show that the fit has similar performance, regarding $-\mathrm{Im}\,\mathcal{A}$ on the real energy axis, as those indicated in Fig. 18.
It is worth pointing out that although we use the $\mathrm{Im}\,E^{tr}_{\alpha}$ parameter to vary the discretized BCTs [^20] This is different from the BCT rotation in the complex scaling method, in which the angle is a controlling parameter., a thorough study of other $E^{tr}_{\alpha}$ distributions and their impacts on the discretized BCTs could reveal more effective controls and thus provide a better solution to the resonance identification task, which could substitute the extra step of forcefully rotating the BCTs in Fig. 19. The so-called Greedy algorithm Sarkar and Lee (2021b); Quarteroni et al. (2016) could help achieve this goal.
@@ V.3 Emulation for parameterized $H$ and sources
First, we fix the sources and study spectrum emulations. When sampling the training points, $\mathrm{Im}\,E^{tr}_{\alpha}$ is fixed, while $\mathrm{Re}\,E^{tr}_{\alpha}$, $\lambda_{4}$ and/or $\lambda$ are sampled using LHS in the corresponding space.
If the systems at the emulation and training points have the same thresholds, as done in Sec. V.3.1, the emulated spectra follow the characteristics seen in Sec. V.2. However, when we vary $\lambda$ and thus the particle-dimer threshold in Sec. V.3.2, the emulated spectra are somewhat distorted in the region away from the threshold(s) at the emulation point. In both sections, $-\mathrm{Im}\,\mathcal{A}$ is accurately reproduced in the emulations, even though system behavior varies dramatically in the parameter space.
Sec. V.3.3 studies the emulations of particle-dimer scattering amplitudes, which require emulations in the source parameters. One source parameter is the scattering energy between the particle and dimer ($E_{rel}$) in the incoming channel. Moreover, the sources depend on $\lambda$ and $\lambda_{4}$. Therefore, the emulation parameter space is five-dimensional in $(\mathrm{Re}\,E,\mathrm{Im}\,E,E_{rel},\lambda,\lambda_{4})$. Again, when sampling the training calculations using LHS, $\mathrm{Im}\,E^{tr}_{\alpha}$ is fixed.
By examining the structure of the sources in Eqs. (55) and (57), we see that $\lambda_{4}$ is an affine parameter, but $E_{rel}$ and $\lambda$ are not—$\lambda$ is affine in $H$. Besides the affine $\lambda$ dependence in $V_{1}$ inside the sources, the dimer (as a subsystem) bound state varies with $\lambda$ in a nonlinear way, which, however, can be emulated using RBM-based bound state emulators König et al. (2020); Ekström and Hagen (2019); Duguet et al. (2024); Drischler et al. (2023); Melendez et al. (2022). The $E_{rel}$ dependence could be emulated using data-driven methods or approximated by affine structures. These extra considerations will be investigated in the future.
Here, we study the accuracy of the scattering amplitude emulations. By probing individual cases and collecting a large testing sample, we see good emulation performance and its variations against the two-body binding energy.
@@ V.3.1 Emulation in $(\mathrm{Re}\,E,\mathrm{Im}\,E,\lambda_{4})$ with fixed sources
Here, the sources are fixed as in Sec. V.2. $\lambda$ is fixed in $H$ so that the dimer binding energy $B_{2}$ and thus the particle-dimer threshold are fixed. $\mathrm{Im}\,E^{tr}_{\alpha}$ is fixed to either $3$ or $10$ MeV, when we sample the training points in $(\mathrm{Re}\,E,\lambda_{4})$ space. In all the results presented here, we sampled $N_{b}=60$ training points, but as will be seen in the results, $N_{\mathrm{eff}}$s are smaller.
@@
Figure 20: Emulations in $(\mathrm{Re}\,E,\mathrm{Im}\,E\lambda_{4})$ space, but with sources fixed as in Fig. 15, and the $B_{2}=10$$\lambda$ is fixed correspondingly. Each row shows the same information about emulation as in Fig. 15, but the rows differ by the $\lambda_{4}$ value at the emulation point.
@@
Figure 20 follows the presentation, for example, in Fig. 15, except the results at different $\lambda_{4}$ values are stacked together vertically. Here $\lambda$ is tuned so that $B_{2}=10$ MeV. The $\mathrm{Im}\,E^{tr}_{\alpha}$ and $\lambda_{4}$ values are shown in the title of each panel.
@@
The compressed spectra from emulations have similar essential characteristics as those in Sec. V.2, including the physical states, the discretized BCTs, and the thresholds. However, a significant difference emerges: the $N_{\mathrm{eff}}$ is significantly bigger than the corresponding value in Sec. V.2, and in terms of spectra, the two BCTs merge at $\mathrm{Im}\,E\sim-50$ MeV, a location much deeper than the merging point ($\mathrm{Im}\,E\sim-10$ MeV) in Fig. 15. This is also expected since the solution subspace’s dimension increases with the parameter space’s dimension. As a result, the differences observed here are also seen in the later plots.
@@
The results extracted using the AAA methods are also shown. Since the AAA algorithm only works for univariate functions, it is not directly feasible to fit it to the training data in a multiple-dimensional parameter space. Therefore, we apply the AAA method at each row in the plot to fit it to our emulation results at the fixed $\lambda_{4}$ but various $E^{tr}_{\alpha}$ values. Essentially, AAA is used here to analytically continue the emulation results in the complex $E$ plane. Consequently, spectra and $N_{\mathrm{eff}}$ from this procedure (still labeled as “3A”) are more similar to those in Sec. V.2. The difference between the emulation and the AAA results suggests that the analytical continuation in the complex plane inside the current emulator differs from the (near)-optimal rational approximation.
@@
As shown in the 3rd and 4th columns, emulators and the AAA procedure reproduce the exact $-\mathrm{Im}\,\mathcal{A}$ very well on the real axis. Again, reducing $\mathrm{Im}\,E^{tr}_{\alpha}$ is needed for improving the near-threshold behavior. Note the errors in the 2nd and 3rd rows are enhanced due to the near-threshold three-body bound states, as seen in Fig. 16.
@@
We note that the plotted $\lambda_{4}$ values are chosen to track the movements of the two three-body bound states with varying $\lambda_{4}$, which are reproduced by the emulators very accurately. Therefore, although our primary interest is in continuum physics, the emulators can also be used to compute the discrete bound states. For that, Sec. IV.2 suggest using a different training point set up to improve the bound state emulations further.
@@
Fig. 21 shows the same results for a smaller $B_{2}$ value ($B_{2}=2$ MeV). The emulator spectra are similar to those in the $B_{2}=10$ MeV case in Fig. 20 but again different from those in Fig. 16. The AAA results, however, only show one discretized BCT, corresponding to the particle-dimer threshold. Remarkably, the emulators and the AAA procedure can reproduce the changes of the spectra as well as the observables $\mathcal{A}$ on the real energy axis, although the results change quite dramatically in the plots.
@@
Note that a bound state could be extremely close to the particle-dimer threshold in some plotted cases. There are no resonances here, but there is a strong signal for virtual states, for example, in the last row of Fig. 21, based on the near-threshold behavior of $-\mathrm{Im}\,\mathcal{A}$ on the real axis. We compute the particle-dimer scattering phase shifts for those $\lambda_{4}$ values. Fig. 22 suggests that the system has virtual states for $\lambda_{4}=-0.2$ and $0.2$. A technical note: the sources in the phase shift scattering vary with scattering energy, which differs from the emulations performed here. However, such a difference is irrelevant when identifying virtual states. Also note that with the BCTs as given, the virtual states are still on the 2nd Riemann sheet and thus can not be identified as an eigenstate; however, the phase shifts can be emulated, as shown in Sec. V.3.3.
@@
We further reexamine the case without a bound dimer in the current situation. The results are shown in Fig. 23, for which $\lambda=-0.7$ is fixed. In all the shown cases, there is a near-threshold broad three-body resonance. Both emulators and AAA with $\mathrm{Im}\,E^{tr}=3$ MeV can reproduce the resonance state, but separating the state from the DC states turns tricky. As discussed around Fig. 17, smaller $\mathrm{Im}\,E^{tr}_{\alpha}$ is needed to increase the density of the DC states to sharpen the separation.
For each of four $\lambda_{4}$ values, the $-\mathrm{Im}\,\mathcal{A}$ on the real axis has a peak feature near the three-particle threshold. It isn’t easy to extract the resonance information based on $-\mathrm{Im}\,\mathcal{A}$ unless the threshold behavior can be separated from the resonance, which again asks for the high DC state density. However, both emulators and the AAA procedure reproduce the function on the real axis.
@@ V.3.2 Emulation in $(\mathrm{Re}\,E,\mathrm{Im}\,E,\lambda,\lambda_{4})$ with fixed sources
We expand the emulations in Sec. V.3.1 by varying both $\lambda$ and $\lambda_{4}$ in $H$ but still using the same sources as in Sec. V.2. $\mathrm{Im}\,E^{tr}_{\alpha}$ is fixed to either $3$ or $10$ MeV, when we sample the training points in $(\mathrm{Re}\,E,\lambda,\lambda_{4})$ space. In all the results presented here, we sampled $N_{b}=60$ training points. Again, $N_{\mathrm{eff}}$ is smaller. The results based on the AAA algorithm are not shown, as the results are qualitatively similar to the AAA results in Sec. V.3.1.
@@
Figure 25: Emulation in the $(\mathrm{Re}\,E,\mathrm{Im}\,E,\lambda,\lambda_{4})$ space. The emulated $-\mathrm{Im}\,\mathcal{A}$ on the real energy axis is compared to the exact results for every $\lambda_{4}$ and $\lambda$ combination whose spectrum is shown in previous Fig. 24. The emulator errors are also shown explicitly.
Fig. 24 shows emulated spectra at different $(\lambda_{4},\lambda)$ values (shown in each panel title). The spectra from $\mathrm{Im}\,E^{tr}_{\alpha}=3$ and $10$ MeV emulators are contrasted in each panel. From left to right, $\lambda$ increases from $-2$ to $0.2$. Accordingly, the dimer becomes less bound in the first two columns and then becomes unbound in the last two; the particle-dimer threshold moves to the right and eventually merges with the three-particle threshold. The three-body bound states also become less bound and could even disappear. From top to bottom, $\lambda_{4}$ increase from $-0.2$ to $0.2$. Generally, the three-body bound states become less bound and eventually disappear.
The emulators reproduce the noted variations of the spectrum properties well. However, in the 3rd column, where a near-threshold resonance exists—to the left of the threshold, it is still on the 2nd Riemann sheet according to the discretized BCTs and thus does not show up in the emulator spectra (see one of the two insets). We then perform a second analysis in each panel, which is the same as the analysis done in Fig. 19, to “rotate” the BCTs even further away from the real axis ($\theta_{\mathrm{BCT}}=3\pi/4$). As shown in the other inset in each panel, the resonance state emerges and agrees well with the exact result.
In parallel to Fig. 24, Fig. 25 shows the corresponding $-\mathrm{Im}\,\mathcal{A}$ emulations on the real energy axis and the emulations errors. These tests show good emulation capabilities at different parameters, although $-\mathrm{Im}\,\mathcal{A}$ varies significantly in magnitude and threshold behavior. The emulation errors behave as before: when reducing $\mathrm{Im}\,E^{tr}$, the errors around the thresholds generally decrease; however, they would persist if there is a near-threshold three-body bound state (see the 2nd column). For the latter, we could eliminate the bound state contribution in $\mathcal{A}$ by projecting out the bound state in the sources, as applied in the LIT studies Efros et al. (2007). Also, note that the two emulators perform almost the same at higher energies.
Figure 26: Emulation in the $(E,\lambda,\lambda_{4})$ space. Here, we test the spectrum emulations by fixing $\lambda_{4}=0$ and varying $\lambda$ in $[-1,-0.7]$. The corresponding trajectory of the emulated resonance in the complex $E$ plane is shown for both emulators. The exact results are also plotted for comparison.
@@
A particular application of the spectrum emulation is quickly mapping out the bound and resonance states’ dependence on the $H$ parameters. To demonstrate that, each panel in Fig. 26 plots an emulated trajectory of a physical state in the complex plane when we vary $\lambda$ from $-1$ to $-0.7$ with $\lambda_{4}=0$, and compares that with the exact results.
@@
According to the exact results, the three-body bound state becomes increasingly less bound, becoming a resonance at $\lambda\sim-0.8$. It continues deeper into the complex plan in the 4th quadrant. These resonances are incredibly close to the three-particle threshold.
@@
By comparing emulations with the exact results, we see bound state emulations are generally better than resonance state emulations. The resonance emulation error increases when the state moves further from the real axis. This is consistent with the increase of extrapolation errors when moving away from the training points in the complex plane. For reader’s information, Fig. 27 shows the emulated spectra, from which we identify the resonance states plotted in Fig. 26.
V.3.3 Emulation in $(\mathrm{Re}\,E,\mathrm{Im}\,E,E_{rel},\lambda,\lambda_{4})$

### v2

@@ V.2 Emulation in ($\mathrm{Re}\,E$, $\mathrm{Im}\,E$) with fixed $H$ and sources
[…]
Fig. 16 shows a set of results with $\theta_{\text{BCT}}=0.75\pi$. In the first two panels, due to the different horizontal scales on the left and right sides of the plots, the discretized BCTs seem bent, but they are on straight lines. With this extra step and a large $\theta_{\text{BCT}}$, the resonances, extracted based on the training data and partial information from the previous emulators, now agree with the exact result. This also shows that the resonance can be identified as an isolated pole and, thus, part of the eigenstates. The 3rd and the 4th panels show that the fit has similar performance, regarding $\mathrm{Im}\,\mathcal{A}$ on the real energy axis, as those indicated in Fig. 15.
It is worth pointing out that although I use the $\mathrm{Im}\,E^{\mathrm{tr}}_{\alpha}$ parameter to vary the discretized BCTs [^16] This is different from the BCT rotation in the complex scaling method, in which the angle is a controlling parameter., a thorough study of other $E^{\mathrm{tr}}_{\alpha}$ distributions and their impacts on the discretized BCTs could reveal more effective controls and thus provide a better solution to the resonance identification task, which could substitute the extra step of forcefully rotating the BCTs in Fig. 16. The so-called greedy algorithm [107, 20] could help achieve this goal.
@@ V.3 Emulation for parameterized $H$ and sources
First, I fix the sources and study spectrum emulations. When sampling the training points, $\mathrm{Im}\,E^{\mathrm{tr}}_{\alpha}$ is fixed, while $\mathrm{Re}\,E^{\mathrm{tr}}_{\alpha}$, $\lambda_{4}$ and/or $\lambda$ are sampled using LHS in the corresponding space.
If the systems at the emulation and training points have the same thresholds, as done in Sec. V.3.1, the emulated spectra follow the characteristics seen in Sec. V.2. However, when I vary $\lambda$ and thus the particle-dimer threshold in Sec. V.3.2, the emulated spectra are somewhat distorted in the region away from the threshold(s) at the emulation point. In both sections, $\mathrm{Im}\,\mathcal{A}$ is accurately reproduced in the emulations, even though system behavior varies dramatically in the parameter space.
Sec. V.3.3 studies the emulations of particle-dimer scattering amplitudes, which require emulations in the source parameters. One source parameter is the scattering energy between the particle and dimer ($E_{\mathrm{rel}}$) in the incoming channel. Moreover, the sources depend on $\lambda$ and $\lambda_{4}$. Therefore, the emulation parameter space is five-dimensional in $(\mathrm{Re}\,E,\mathrm{Im}\,E,E_{\mathrm{rel}},\lambda,\lambda_{4})$. Again, when sampling the training calculations using LHS, $\mathrm{Im}\,E^{\mathrm{tr}}_{\alpha}$ is fixed.
[…]
For now, I study the accuracy of the scattering amplitude emulations. By probing individual cases and collecting a large testing sample, one can see good emulation performance and its variations against the two-body binding energy.
@@ V.3.1 Emulation in $(\mathrm{Re}\,E,\mathrm{Im}\,E,\lambda_{4})$ with fixed sources
Here, the sources are fixed as in Sec. V.2. $\lambda$ is fixed in $H$ so that the dimer binding energy $B_{2}$ and thus the particle-dimer threshold are fixed. $\mathrm{Im}\,E^{\mathrm{tr}}_{\alpha}$ is fixed to either $3$ or $10$ MeV, when sampling the training points in $(\mathrm{Re}\,E,\lambda_{4})$ space. In all the results presented here, I sampled $N_{b}=60$ training points, but as will be seen in the results, $N_{\mathrm{eff}}$s are smaller.
@@
Figure 17: Emulations in the $(\mathrm{Re}\,E,\mathrm{Im}\,E,\lambda_{4})$ space. The sources are the same as in Fig. 12. I choose $B_{2}=10$, which fixes $\lambda$ accordingly. Each row shows the same information as in Fig. 12, but the rows differ by the $\lambda_{4}$ value at the emulation point.
@@
Figure 17 follows the presentation, for example, in Fig. 12, except that the results at different $\lambda_{4}$ values are stacked together vertically (more results in the SM [110]). Here, $\lambda$ is tuned so that $B_{2}=10$ MeV. The $\mathrm{Im}\,E^{\mathrm{tr}}_{\alpha}$ and $\lambda_{4}$ values are shown in the title of each panel.
@@
The compressed spectra from emulations have similar essential characteristics as those in Sec. V.2, including the physical states, the discretized BCTs, and the thresholds. However, a significant difference emerges: the value of $N_{\mathrm{eff}}$ is significantly bigger than the value in Sec. V.2, and in terms of spectra, the two BCTs merge at $\mathrm{Im}\,E\sim-50$ MeV, a location much deeper than the merging point ($\mathrm{Im}\,E\sim-10$ MeV) in Fig. 12. This is also expected since the solution subspace’s dimension increases with the parameter space’s dimension. For the same reason, the differences observed here are also seen in the later plots.
@@
The results extracted using the AAA algorithm are also plotted. Since the algorithm only works for univariate functions, it is not feasible to fit the rational form in Eq. (8) to the training data in a multiple-dimensional parameter space. Instead, I fit it to the emulation results at the fixed $\lambda_{4}$ but different $E^{\mathrm{tr}}_{\alpha}$ values. Essentially, AAA is used to analytically continue these emulation results in the complex $E$ plane. Thus, spectra and $N_{\mathrm{eff}}$ from this procedure (still labeled as “3A”) are more similar to those in Sec. V.2 than the CERPE results are. This difference also suggests that the analytical continuation in CEREP is not the same as the (near-)optimal rational approximation.
@@
In addition, emulators and the AAA procedure reproduce the exact $\mathrm{Im}\,\mathcal{A}$ very well on the real axis, as shown in the 3rd and 4th columns. Again, reducing $\mathrm{Im}\,E^{\mathrm{tr}}_{\alpha}$ is needed to better predict the near-threshold behavior. Note the errors in the 2nd and 3rd rows are enhanced due to the near-threshold three-body bound states, as seen in Fig. 13.
@@
In the plot, $\lambda_{4}$ (including those in the SM [110]) is varied to trace the movements of the two three-body bound states, which are accurately reproduced by the emulators. Therefore, although my primary interest is in continuum physics, these emulators can also be used to compute the discrete bound states. For that, Sec. IV.2 suggest a better training-point setup.
@@
Fig. 18 shows the same results for a smaller $B_{2}$ value ($B_{2}=2$ MeV) with its extended version in the SM [110]. The emulator spectra are similar to those in the $B_{2}=10$ MeV case in Fig. 17 but again different from those in Fig. 13. The AAA results, however, only show one discretized BCT, corresponding to the particle-dimer threshold. Remarkably, the emulators and the AAA procedure can reproduce the spectra and $\mathcal{A}$ on the real energy axis, although they change dramatically when varying $\lambda_{4}$.
@@
Note that a bound state could be extremely close to the particle-dimer threshold in some plotted cases. There are no resonances here, but virtual states could exist, for example, in the first row of Fig. 18, based on the near-threshold behavior of $\mathrm{Im}\,\mathcal{A}$ on the real axis. The particle-dimer scattering phase shifts in Fig. 19 also points to the presence of virtual states for $\lambda_{4}=-0.2$ and $0.2$ (see the SM [110] for the $\lambda_{4}=0.2$ case). It is worth mentioning that the sources in the phase shift scattering vary with $E_{\mathrm{rel}}$, while they are fixed in the emulations performed here. However, such a difference is irrelevant when identifying virtual states. Also note that with the BCTs as given, the virtual states are still on the 2nd Riemann sheet and thus can not be identified as an eigenstate; however, the phase shifts can be emulated, as shown in Sec. V.3.3.
@@
I further reexamine the case without a bound dimer in the current situation. Fig. 20 (and its extension in the SM [110]) shows the examples with $\lambda=-0.7$ and varying $\lambda_{4}$. For each parameter set, there is a near-threshold broad three-body resonance. Both emulators and AAA with $\mathrm{Im}\,E^{\mathrm{tr}}=3$ MeV can reproduce the resonance state, but separating the state from the DC states turns tricky. As discussed around Fig. 14, smaller $\mathrm{Im}\,E^{\mathrm{tr}}_{\alpha}$ is needed to increase the density of the DC states to sharpen the separation.
As related, the peak feature of $\mathrm{Im}\,\mathcal{A}$ on the real axis also sits extremely close to the three-particle threshold. As the result, it is difficult to extract the resonance information based on $\mathrm{Im}\,\mathcal{A}$ unless the threshold behavior can be separated from the resonance. This again asks for the high DC state density. Do note that both emulators and the AAA procedure reproduce $\mathrm{Im}\,\mathcal{A}$ on the real axis.
@@ V.3.2 Emulation in $(\mathrm{Re}\,E,\mathrm{Im}\,E,\lambda,\lambda_{4})$ with fixed sources
I expand the emulations in Sec. V.3.1 by varying both $\lambda$ and $\lambda_{4}$ in $H$ but still using the same sources as in Sec. V.2. $\mathrm{Im}\,E^{\mathrm{tr}}_{\alpha}$ is fixed to either $3$ or $10$ MeV, when I sample the training points in the $(\mathrm{Re}\,E,\lambda,\lambda_{4})$ space. For all the results presented here, a $N_{b}=60$ number of training points are sampled. But, as before, $N_{\mathrm{eff}}$ is smaller. The results based on the AAA algorithm are not shown, as the results are qualitatively similar to the AAA results in Sec. V.3.1.
@@
Figure 22: Emulation in the $(\mathrm{Re}\,E,\mathrm{Im}\,E,\lambda,\lambda_{4})$ space. The emulated $\mathrm{Im}\,\mathcal{A}$ on the real energy axis is compared to the exact results for every $\lambda_{4}$ and $\lambda$ combination whose spectrum is shown in previous Fig. 21. The emulator errors are also shown explicitly.
Figure 21 and its extension in the SM [110] plot emulated spectra at different $(\lambda_{4},\lambda)$ values (provided in each panel title). The spectra from $\mathrm{Im}\,E^{\mathrm{tr}}_{\alpha}=3$ and $10$ MeV emulators are contrasted in each panel. From left to right, $\lambda$ increases from 2-2 to $0.2$. Accordingly, the dimer becomes less bound in the first two columns and then becomes unbound in the last two; the particle-dimer threshold moves to the right and eventually merges with the three-particle threshold. See the SM [110] for more results with increasing $\lambda_{4}$, where the three-body bound states become less bound and eventually disappear.
The emulators reproduce the noted variations of the spectrum properties well. However, in the 3rd column, where a near-threshold resonance exists—to the left of the threshold, it is still on the 2nd Riemann sheet according to the discretized BCTs and thus does not show up in the emulator spectra (see one of the two insets). I then perform a second analysis in each panel, which is the same as the analysis done in Fig. 16, to “rotate” the BCTs even further away from the real axis ($\theta_{\mathrm{BCT}}=3\pi/4$). As shown in the other inset in each panel, the resonance state emerges and agrees well with the exact result.
In parallel to Fig. 21, Fig. 22 shows the corresponding $\mathrm{Im}\,\mathcal{A}$ emulations on the real energy axis and the emulations errors. These tests (including the extended figure in the SM [110]) show good emulation capabilities at different parameters, although $\mathrm{Im}\,\mathcal{A}$ varies significantly in terms of magnitude and threshold behavior. The emulation errors behave as before: when reducing $\mathrm{Im}\,E^{\mathrm{tr}}$, the errors around the thresholds generally decrease; however, they would persist if there is a near-threshold three-body bound state (see the 2nd column). For the latter, one could eliminate the bound state contribution in $\mathcal{A}$ by projecting out the bound state in the sources, as applied in the LIT studies [67]. Also, note that the two emulators perform almost the same at higher energies.
Figure 23: Emulation in the $(E,\lambda,\lambda_{4})$ space. Here, I test the spectrum emulations by fixing $\lambda_{4}=0$ and varying $\lambda$ in [1,0.7][-1,-0.7]. The corresponding trajectory of the emulated resonance in the complex $E$ plane is shown for both emulators. The exact results are also plotted for comparison.
@@
A particular application of the spectrum emulation is quickly mapping out the bound and resonance states’ dependence on the $H$ parameters. To demonstrate that, each panel in Fig. 23 plots an emulated trajectory of a physical state in the complex plane when $\lambda$ is varied from 1-1 to 0.7-0.7 with $\lambda_{4}=0$.
@@
According to the exact results shown there, the three-body bound state becomes increasingly less bound, turns into a resonance at $\lambda\sim-0.8$, and moves deeper into the complex plan in the 4th quadrant. These resonances are incredibly close to the three-particle threshold.
@@
One can see that the bound state emulations are generally better than the resonance state emulations. The resonance emulation error increases when the state moves further away from the real axis. This is consistent with the increase of extrapolation errors when moving away from the training points in the complex plane. For reader’s information, Fig. 24 shows the emulated spectra, from which I identify the resonance states plotted in Fig. 23.
V.3.3 Emulation in $(\mathrm{Re}\,E,\mathrm{Im}\,E,E_{\mathrm{rel}},\lambda,\lambda_{4})$